# GARLIC: GAussian Representation LearnIng for spaCe partitioning

### v1

@@ A.2 Implementation details
A.3 Additional Ablation Studies
[…]
(a) Loss terms
(b) Split & clone impact
(c) ReLU parameters $\tau$
(d) High-recall
[…]
Architectural components.
In Fig. 5(5(a)), the contribution of each loss term to the optimization process is analyzed. The results indicate that the covariance loss $\mathcal{L}_{\text{cov}}$ produces the most significant performance improvement, while the anchor loss $\mathcal{L}_{\text{anchor}}$ functions as an effective regularizer, grounding each Gaussian by aligning it with its corresponding point distribution. Fig. 5(5(b)) demonstrates the effects of our adaptive refinement operations, where the split and clone mechanisms improve retrieval performance when used together, each contributing by providing complementary benefits to the quality of representation. Fig. 5(5(c)), addresses the parameter $\tau$ within $\mathcal{L}_{div}$, demonstrating that $\tau=3$ is the optimal choice, particularly in scenarios involving a limited number of probes. Values surpassing those presented in our study ($\tau>3$) result in less optimal outcomes. For instance, $\tau=4$ demonstrates a Recall10@10 value of 0.670.67~{}0.670.67 when evaluated with $65000$ probes.
A way to address the limitation associated with high-recall scenarios, as discussed in the limitations paragraph of the main paper, is show in Fig. 5(5(d)). Specifically, it involves assessing all Gaussian candidates that meet the condition $d_{M}(x,g)\leq\tau$, rather than solely the one that minimizes the Mahalanobis distance. This approach is also semantically justified, considering that multiple Gaussians may occupy the same region. Consequently, a hybrid approach can be used, utilizing the $\arg\min$ strategy for scenarios with a low number of probes and employing all candidates that satisfy condition $d_{M}(x,g)\leq\tau$ for high-recall cases, thus effectively addressing the limitation.
(a) Initialization
(b) Split threshold $\gamma$
(c) Clone threshold $e$
(d) Dimensionality effect
[…]
Table 4: Effect of anisotropy on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better. As anisotropy decreases, performance degrades due to excessive probe usage.
Configuration Performance $\uparrow$
Covariance structure
   Anisotropic 10.43
   Diagonal 11.04
   Isotropic 10.63
Validation of hyperparameters.
Fig. 6 provides a systematic evaluation of our key hyperparameters. In Fig. 6(6(a)), we investigate initialization strategies for Gaussian centers, revealing that K-Means initialization consistently outperforms alternatives. Although farthest point sampling (FPS) achieves comparable recall, it requires approximately 50% more probes, and catching up to K-Means requires more training time. In Fig. 6(6(b)), the split threshold $\gamma$, controlling the density condition for Gaussian subdivision, is examined. The optimal values are $\gamma=0.005$ and $\gamma=0.01$, though $\gamma=0.005$ requires more time due to more frequent splits. As for the clone threshold $e$, depicted in Fig. 6(6(c)), which determines the outer shell for cloning, it remains stable across different values, with $e=2.2$ being preferable in situations requiring high recall. Finally, Fig. 6(6(d)) examines dimensionality effects using Fashion-MNIST data resized to various dimensions. The results confirm that GARLIC maintains strong performance across a wide range of dimensionalities, highlighting its robustness to the feature space complexity.
Covariance structure vs performance.
We conduct a targeted study to assess the impact of different covariance configurations-namely, full (anisotropic), diagonal, and isotropic-baffect the performance of GARLIC. As shown in Table 4, reducing the Gaussian expressiveness from full to diagonal and then to isotropic leads to a notable decline in average Recall10@10 per probe. While diagonal and isotropic configurations reduce both the parameter count and computation per Gaussian to a linear level, $\mathcal{O}(k\cdot d)$, they lead to aggressive pruning and an increased number of probes to make up for the representational loss. This suggests that space complexity cannot be drastically reduced without harming retrieval quality because simpler Gaussian parameterizations result in degraded locality and coverage.
A.4 Qualitative Results
Coverage and reconstruction diagnostics.
Curvature-based diagnostics.
Fig. 8 provides four views exploring the relationship between curvature and structural properties of the Gaussian assignments across datasets, to examine wheather the learned model is informative and geometrically consistent with the data $x\in X$.
Fig. 9 demonstrates the capability of GARLIC to capture the intrinsic geometric structure of high-dimensional data with locally varying dimensionality. The anisotropy histograms (first row) reveal how Gaussians adapt to regions of different local dimensionality, with $\lambda_{\text{max}}/\lambda_{\text{min}}$ ratios ranging from nearly isotropic to highly stretched configurations across all datasets. Rather than simply partitioning space uniformly, Gaussians adapt their shapes to the underlying manifold structure, as confirmed by the diagonality histograms (third row) and effective dimension measurements (fourth row). This representation allows us to model stratified data where different intrinsic dimensionalities coexist, allowing the data to become pancake-like for surface regions, needle-like for curve regions, and ball-like for volumetric areas. Unlike traditional space partitioning methods, GARLIC models the underlying data distribution probabilistically, not just approximating distances for retrieval. The explained variance surfaces (bottom row) show that Gaussians efficiently capture the relevant dimensions at each location, enabling estimation of true manifold distances rather than just Euclidean distances to samples. This provides more semantically meaningful results in regions where the intrinsic dimensionality is lower than the ambient space. This completes the supplementary material.

### v2

@@ A.2 Implementation details
A.3 Complexity Analysis
Index build complexity.
Let $I$ be the number of optimization steps and $K^{\prime}$ the initial number of Gaussians. We denote by $S$, $C$, and $P$ the number of splits, clones, and pruned Gaussians, respectively, and define the final number of Gaussians as $K=K^{\prime}+S+C-P$. Let $|B_{g}|$ be the average cell size, $c$ the number of K-Means clusters used during splitting, and $k^{\prime}$ the number of candidate points sampled per cloning operation.
For the initialization, since we use K-Means++ on $K^{\prime}$ total cluster centers, we need $\mathcal{O}(K^{\prime}\cdot d\cdot|X|)$ time. For the optimization part, we need to perform a total number of $I$ iterations of full Mahalanobis-based point-to-Gaussian assignment, thus a total of $\mathcal{O}(I\cdot|X|\cdot K\cdot d^{2})$ worst-case time. Separate from the optimization, we analyze the split, clone and prune operations that are not applied on every iteration of the optimization. (i) The split operation runs DBSCAN or K-Means (with $c$ clusters), thus for a total of $S$ such operations we would need $\mathcal{O}(S\cdot(|B_{g}|\cdot d\cdot c+d^{2}))$; (ii) the clone operation locates the subset of points outside the Gaussian’s boundary (between $\tau\cdot\sigma$ and $(1+\epsilon)\cdot\tau\cdot\sigma$) for which it identifies new local modes, thus for a total of $C$ operations, it leads to $\mathcal{O}(C\cdot(k^{\prime 2}\cdot d+d^{2}))$; and (iii) the prune operation simply removes low-cardinality Gaussians and reassigns the points to the nearest active Gaussian, which takes $\mathcal{O}(P\cdot|B_{g}|\cdot d)$ time. For the quantization of each Gaussian, PCA is performed on all points inside the Gaussian, which projects the data into reduced local bases. In total, for the quantization we need $\mathcal{O}(K^{\prime}\cdot|B_{g}|\cdot d^{2}+|X|\cdot d\cdot r)$ time. From all the terms described, the optimization term dominates.
Query complexity.
Space complexity.
Let $K$ be the number of Gaussians, $d$ the data dimension, and $N=|X|$ the dataset size. The model stores mean vectors $\boldsymbol{\mu}\in\mathbb{R}^{K\times d}$, Cholesky parameters $\mathbf{L}\in\mathbb{R}^{K\times d\times d}$, and cells storing point indices, requiring $\mathcal{O}(N)$ space. Thus, the total space complexity is:
$\mathcal{O}\left(K\cdot(d^{2}+d)+N\right)$